OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (xguardians):

What method are you using? Substitution?

OpenStudy (anonymous):

take y seperate from 2nd equation and put it into 1st u will have value of x

OpenStudy (xguardians):

Okay, one second.

OpenStudy (anonymous):

u can hv another method .. multiplying 2nd equation with 2 and than add the resultant equaton and 1st one .. u will have only x in answer

OpenStudy (anonymous):

I don't really understand what you mean, can you show me the steps?

OpenStudy (xguardians):

y= -3x + 6 Input it in. x - 2(-3x + 6) = 2 x +6x - 12 = 2 7x = 14 x = 2 Then replug that back into one of the equations... 3(2) + y = 6 y + 6 = 6 y = 0

OpenStudy (xguardians):

1. So first you isolate a variable in one of them. 2. Input that into the other equation. 3. Solve for a variable 4. Replace that variables value into either of equations 5. Solve for other variable 6. Finish.

OpenStudy (anonymous):

oh, so you change the equation to equal y then use that to solve for x. thanks :)